# Learning to Generate Novel Scientific Directions with Contextualized Literature-based Discovery

### 22 Jan 2024 v4

@@ 4 Experiments
4.1 Dataset
We construct our dataset from 67,408 ACL Anthology papers from S2ORC Lo et al. (2020). Given a title and the corresponding abstract from a document $d$, we first apply state-of-the-art scientific IE system PL-Marker Ye et al. (2022) to extract nodes $v\in V$ and relations which are edges $e\in E$ between nodes. Next, we use SciCo Cattan et al. (2021) to obtain coreference links. Then, we use ScispaCy Neumann et al. (2019) to replace abbreviations with a more informative long form. Finally, we perform scientific sentence classification Cohan et al. (2019) to find background context $\mathcal{B}$. We also collect paper metadata, including the citation network $\mathcal{G}_{c}$. We split our dataset temporally (train/dev/test correspond to papers from years $<$2021 / 2021 / 2022 respectively). Table 1 shows data statistics. More details are in Appendix A.
Quality of IE Preprocessing
During preprocessing, we only keep high-confidence outputs from IE models to reduce errors. We observe this removes many of the noisy cases. To validate this, we manually evaluate precision of each preprocessing step on a random sample of papers and observe that all steps yield high precision (%91-100%) except relation extraction (%65); in total, the rate of instances passing all steps was 79.7%. See Table 6 in Appendix and data limitations in §7.
4.2 Idea Sentence Generation Evaluation
We begin with automatic evaluation, followed by human evaluation in §5. We use standard evaluation metrics, including ROUGE Lin (2004), BERTScore Zhang* et al. (2020) and BARTScore Yuan et al. (2021), to check the similarity between ground truth and generated output.
Subset Challenging Gold
GPT3.5ZS
0.136
0.624
0.147
0.631
GPT3.5FS
0.189
0.641
0.210
0.655
GPT3.5Retr
0.185
0.642
0.199
0.653
GPT4ZS
0.120
0.581
0.130
0.583
GPT4FS
0.143
0.618
0.151
0.624
T5
0.223
${}^{\dagger}$
0.246
0.685
[…]
GPT3.5FS+SN
0.183
0.640
0.191
0.649
GPT3.5FS+KG
0.187
0.642
0.201
0.653
GPT3.5FS+CT
0.169
0.641
0.182
0.649
GPT4FS+KG
0.143
0.619
0.152
0.626
T5+CL
${}^{\dagger}$
${}^{\dagger}$
${}^{\dagger}$
${}^{\dagger}$
T5+SN+CL ${}^{\dagger}$
${}^{\dagger}$
${}^{\dagger}$
${}^{\dagger}$
T5+KG+CL
${}^{\dagger}$
0.669
0.248
${}^{\dagger}$
T5+CT+CL
${}^{\dagger}$
${}^{\dagger}$
${}^{\dagger}$
${}^{\dagger}$
Table 2: Automatic evaluation results on idea sentence generation for the challenging and gold subsets. CL is a model with in-context contrastive augmentation. SN is a model with semantic neighbors. KG is a model with KG neighbors. CT is a model with citation neighbors. R-L denotes ROUGE-L. BERT denotes BERTscore with SciBERT as its encoder. ${}^{\dagger}$ indicates that differences between models are not statistically significant ($p\leq 0.05$) when compared to each other but are still significant when compared to the other models.
We remove test instances where models can trivially use surface-level background information to infer the ground truth. We select test instances with low similarity between background context and corresponding ground truth sentences. To create this subset, we compute the cosine similarity between each instance’s background context and its corresponding ground truth sentence in the test set. We then choose pairs with cosine similarity lower than $0.074$, which amounts to the tenth percentile of pairs with the lowest cosine similarity.
We further annotate this subset to create a gold subset. We manually exclude instances with trivial overlap between the ground truth sentence and background context. We retain only instances where the background context is relevant to the target relation and where the target relation (from which the seed term is taken) is central to the target paper’s idea. We also removed test pairs that have abbreviations or unexplained terms in the background context. Full annotation details are in Appendix G. We obtained 194 instances satisfying those conditions. [^3] We also evaluate the full dataset without any filters, obtaining higher results as expected but with similar overall findings (see Table 9 in the Appendix).
In Table 2, our inspiration-enhanced models achieve superior performance. GPT-FS models are outperformed by all other T5-based models. In addition, GPT-FS (few-shot) approaches are considerably better than GPT-ZS (zero-shot). [^4] We also collected human ratings for GPT4-ZS vs. GPT4-FS using the criteria in §5; GPT4-FS ranked higher in 65% of cases, with the rest mostly tied. We find GPT4 outputs to be much longer than GPT3.5D), explaining why they underperform in automatic metrics but outperform GPT3.5 in human evaluations (§5). Finally, our in-context contrastive examples, which encourage novelty with respect to background context, help models perform better than baselines in most cases. Results from the semantic and citation graphs have better performance than with KG — many test instances with new or rare concepts do not have any KG neighbors; however, the KG can guide GPT4 to compose more concrete ideas (§5 and Table 3). The main takeaways, though, are that our task is challenging even for strong large language models, and that automatic evaluation metrics are highly problematic in our setting (e.g., ROUGE-L shows low variation across the different models partially due to the generated sentences often following certain templates (e.g. “In this paper, we propose a new … for …”); we conduct human evaluations and analysis in §5.
Type Content
Input
seed: class imbalance is done by using; context (abridged): … class imbalance and drift can sometimes be mitigated by resampling the training data to simulate a known target distribution, but what if the target distribution is determined by unknown future events?
[…]
GPT4FS
In that context, we use resampling the training data to address class imbalance, as it helps simulate or compensate for a known target distribution even when dealing with unknown future events.
[…]
GPT4FS + KG
In that context, we use a method called Frustratingly Easy Model Averaging (FEDA) to handle class imbalance, as it combines the outputs of multiple weak learners trained on different subsets of the training data. This approach helps in mitigating the impact of class imbalance by leveraging the predictions of various models, thus providing a more robust and generalized decision. Additionally, FEDA can adapt to temporal concept drift, as it considers the predictions from different models, which might capture diverse aspects.
[…]
Ground Truth
Reframing group-robust algorithms as adaptation algorithms under concept drift, we find that Invariant Risk Minimization and Spectral Decoupling outperform sampling approaches to class imbalance and concept drift.
Table 3: Comparing GPT4FS with GPT4FS+KG and the original idea sentence (texts cut to fit). Overall, our analysis finds the GPT4FS+KG leads to suggestions with improved technical depth. However, both are far from the technical depth of innovation in original human-proposed ideas.
4.3 Idea Node Prediction Evaluation
Adding citation neighbors leads to a performance increase for SciBERT, CSRoBERTa, and T5. Generally, however, results indicate the task’s highly challenging nature, both in modeling and evaluation. We present and discuss results in Appendix E. We also conduct a multi-choice evaluation, to test the models’ ability to distinguish between relevant candidates and distractors (Appendix E.3).
5 Human Evaluation & Error Analysis
We conduct human evaluations of model quality; finally, we also demonstrate the generalization of our approach to the biomedical domain.
Study I
We recruit six volunteer NLP experts with graduate-level education to rate the system. Raters are told to envision an AI paper-writing assistant that suggests new ideas. We randomly select 50 instances (background+seed) from the sentence generation gold subset (§A). Each annotator is given ten instances, each paired with five system outputs including: GPT3.5FS, GPT3.5Retr, GPT3.5FS+CT, GPT3.5FS+KG, GPT4FS, GPT4FS+KG, T5, and T5+SN+CL. Raters are blind to the condition, and system outputs are randomly shuffled across instances. We observe moderately high rater agreement (Table 24 in the Appendix). We ask raters to assess idea quality by taking into account each output’s relevance to the context, novelty, and whether the idea is clear and reasonable (positive ratings are dubbed “helpful” as shorthand corresponding to passing the multiple considerations). We instruct annotators to only provide positive ratings to ideas that are sufficiently different from the input context. In Study I, we ask raters not to anticipate groundbreaking novelty from the system but rather a narrower expectation of quality and utility; in Study II, below, we enrich the analysis to examine ranking between top models and also “raise the bar” and compare to actual scientific ideas from papers. Further details and evaluator guidelines are in Appendix H (sample annotations in Table 22).
Study I: Results
Overall, GPT4FS and GPT4FS+KG outperform other models by a wide margin (see Table 4). Apart from GPT4, T5+SN+CL performs best compared to other baselines, given its stronger prior knowledge of useful background citations. In general, GPT3.5 models performed significantly worse than fine-tuned T5 and its variants, which echoes results in other work in the scientific NLP domain Jimenez Gutierrez et al. (2022). GPT4 outputs are longer than other baselines (Appendix D), which may partially explain higher human preference.
@@ 5 Human Evaluation & Error Analysis
Table 4: Percent (%) of total votes each system output receives from human raters for the idea sentence prediction task. H denotes a helpful output, while U denotes an unhelpful output. “3FS” refers to the GPT3.5FS. “3Rt” refers to the GPT3.5Retr. “4” refers to GPT4FS, and “4+KG” refers to the GPT4FS+KG. “T5+SN” refers to the T5+SN+CL. GPT4FS and GPT4FS+KG are rated much higher. While GPT4FS has a slightly higher rating than the KG variant, a further human study reveals that GPT4FS+KG often leads to more technical depth (§5).
Study II: GPT4 comparisons to real papers
@@
We conduct a follow-up human study of close competitors GPT4FS and GPT4FS+KG with a subset of the annotators. In this study, model outputs are now ranked, unlike the binary classification of helpful/not in Study I. Suggestions are ranked according to the level of technical detail and innovation in comparison to each other—i.e., ranking which of GPT4FS and GPT4FS+KG had a higher degree of technical detail and novelty, or whether they are roughly the same (tied). Finally, outputs are rated versus the ground truth idea, according to whether or not the suggestions were roughly at the same level of technical detail and innovation as the original paper’s idea, or significantly lower.
Study II: Results
@@
Overall, GPT4FS+KG is found to have higher technical detail in 48% of the compared pairs, and found to be less incremental (more novel) in 45% of the pairs. Among the remaining 52%/55% (respectively), the vast majority are ties, indicating that whenever GPT4FS+KG is not favored, it is of roughly the same quality as GPT4FS, but not vice versa. However, the most crucial aspect is comparing the results against the original ground truth idea on the quality of innovation. Here, we find that in 85% of comparisons, the ground truth is considered to have significantly higher technical level and novelty; and in the remaining 15%, the ground truth was ambiguous or lacking additional context from the paper abstract. This points to a major challenge in obtaining high-quality idea generations using existing state-of-the-art models.
Error Analyses
Models often made generic suggestions, woven together with specific details copied directly from the context (e.g., “NLP with ML algorithms and sentiment analysis” for some problem X, or “data augmentation and transfer learning” for Y, or “BERT or RoBERTa” for Z). Our few-shot prompting techniques reduced this behavior but did not fully solve it. GPT4 models, especially, seemed to generate generic descriptions of common steps in NLP workflows (e.g., “Data preprocessing: Clean the text data, remove unnecessary characters, perform tokenization…”). In general, all models tended more often than not to copy and rephrase substantial amounts of content directly from the context. In certain cases, models applied simple logical modifications to the context; e.g., when contexts described problems such as “high latency” or “efficiency limitations”, the suggestions would include phrases such as “low latency” or “highly efficient”. Further in-depth analyses appear in Appendices E.3,F.
Domain Generalization Case Study
@@
Our framework is domain-agnostic and can be applied to other domains by changing the IE system used in the preprocessing procedure. To demonstrate this, we conduct an additional initial experiment in the biochemical domain. We follow a similar data creation procedure as for NLP papers, outlined here briefly (full details are in Appendix I.1). We collect a dataset from PubMed papers, and use PubTator 3 Islamaj et al. (2021); Wei et al. (2022); Luo et al. (2023); Wei et al. (2023); Lai et al. (2023) as an IE system to extract a KG from paper abstracts. We use a sentence classifier trained on annotated abstracts Huang et al. (2020) to select background context. We fine-tune a state-of-the-art biomedical large language model Wu et al. (2023) on our data [^5] The input and sample outputs are in Table 23.. We ask two biochemical domain experts with graduate-level education to evaluate the quality of the results as before, finding them to overall rate 65% of the generated directions positively. Finally, in contrast to NLP-domain experiments, evaluators were more satisfied with the generated outputs compared to the ground truth in terms of technical detail—finding roughly 50% to be on par with the ground truth in that regard. However, this preliminary experiment was meant mainly to demonstrate the generality of our approach, and a more in-depth exploration of the utility and quality of generated outputs is left for future work.

### 17 Feb 2024 v5

@@ 2.2 Automated Training Data Collection
3 SciMON Models
We present a new module to retrieve inspirations as contextual input (§3.1). Then, we describe another module to generate ideas given the context+inspiration (§3.2). Finally, we introduce a new iterative novelty optimization method to further improve idea quality (§3.3).555 Training and hyperparameter details in Appendix B.
3.1 Inspiration Retrieval Module
We take broad inspiration from cognitive aspects of innovation Hope et al. (2023): when researchers generate a new idea, they are grounded in a web of existing concepts and papers bearing on the new idea. We aim to enrich the context of each background by retrieving “inspirations”— pieces of information that can guide hypothesis generation. As illustrated in Figure 2, for a given instance of the SciMON task, our retrieval augmentation can retrieve from three types of sources. Each source uses a different form of query and output.
Semantic Neighbors For a given problem/motivation as input, ideas proposed for related problems in the training set can serve as a guiding reference for generating a new idea. Given the background context $\mathcal{B}$ with a seed term $v$ and problem/motivation $\mathcal{M}$, we construct a base input $b$: a concatenation of $\mathcal{M}$ with a prompt $\mathcal{P}$ belonging to one of two templates: “$v$ is used for $p$” or “$v$ is done by using $p$”, where $p$ is one of Task/Method/Material/Metric. In short, $b:=$ $\mathcal{P}$ $\oplus$ $\mathcal{M}$. For example, in Figure 2, the concatenation is “Knowledge acquisition is done by using Method; Context:…requires plms to integrate information…lifelong manner…”.
We then retrieve inputs from the training set that are semantically related to a new base input $b$, and obtain target sentences $T$ corresponding to each retrieved training input. We extract the target term $t\in\mathcal{T}$ matching the seed term in $b$2.2) as inspiration for input $b$. Simply put, this means we use as inspiration the salient aspect of the solution proposed in $\mathcal{T}$, which we found empirically to help remove noisy/irrelevant information in $\mathcal{T}$. For example, in Figure 2, we find “informative entities are done by using Method context: in this work, we aim at equipping pre-trained language models with structured knowledge.” as similar to the input and use $t=$“linked knowledge graph” as inspiration.
Technically, we first construct a fully connected graph $\mathcal{G}_{S}$ based on the training set where each node is a pair of input text $b_{i}$ and target term $t_{i}$. We define the weight between two nodes $i$ and $j$ as the cosine similarity between $b_{i}$ and $b_{j}$ based on representations from SentenceBERT (Reimers and Gurevych, 2019) (all-mpnet-base-v2). Given $b$, we first insert it into $\mathcal{G}_{S}$ and compute the weights of its connected edges. We then retrieve neighbors input text $\{b_{1},\ldots,b_{k}\}$ from the training set with the largest edge weight, where $k$ is the number of retrieved instances. We consider the corresponding target terms $\{t_{1},\ldots,t_{k}\}$ as semantic inspirations.
KG Neighbors We also explore enriching the context by linking it to a background KG with information on related methods and tasks. Using the same IE process used to extract our training examples (§2.2), we create a global background KG $\mathcal{G}_{B}$ which covers all papers in the corpus $\mathcal{D}_{\mathcal{Y}}$ prior to a given year $\mathcal{Y}$ (i.e., the nodes in $\mathcal{G}_{B}$ correspond to tasks/methods/materials/metrics, and the edges are used-for relations, extracted and normalized from across the entire corpus as described earlier). Then, given a seed term $v$ at query time, we select adjacent nodes $\{n_{1},n_{2},...\}$ from $\mathcal{G}_{B}$ as inspirations. As an example, in Figure 2, the neighbor nodes of “knowledge acquisition” include “collaborative web text annotation editor”, “image matching”, etc., which we select as inspirations.
Citation Neighbors Another notion of contextual relatedness we explore is via citation graph links. Here, given as input background context $\mathcal{B}$, we assume access to the original source document $d$ from which $\mathcal{B}$ was extracted, and consider its cited paper title set $\mathcal{C}_{d}$ as potential candidates. This can be seen as a stronger assumption on information available to the model— assuming a researcher using the model provides relevant candidate documents from which ideas could be pooled. Because the training set only contains papers before year $\mathcal{Y}$, we only select papers $\mathcal{C}_{d\mathcal{Y}}\subseteq\mathcal{C}_{d}$ prior to year $\mathcal{Y}$. We then retrieve the top-$k$ titles with the highest cosine similarity to $d$ from $\mathcal{C}_{d\mathcal{Y}}$ based on their SentenceBERT embeddings as earlier. For instance, in Figure 2, the paper ELLE Qin et al. (2022) cites the paper de Masson d'Autume et al. (2019). Therefore, we choose the title “episodic memory in lifelong language learning” as inspiration information.
3.2 Generation Module
The idea generation module is given retrieved inspirations $i_{1},\ldots,i_{k}$ along with context $\mathcal{M}$ as input.
In-Context Learning We experiment with recent state-of-the-art LLMs, GPT3.5 davinci-003 Ouyang et al. (2022) and GPT4 gpt-4-0314 checkpoint OpenAI (2023). We first ask the model to generate sentences based on the seed term and the context in the zero-shot setting without any in-context examples (GPT3.5ZS, GPT4ZS). We then ask the model to generate sentences in a few-shot setting by prompting randomly chosen pairs of input and output from the training set (GPT3.5FS, GPT4FS). Inspired by Liu et al. (2022), we further employ a few-shot setting using semantically similar examples. Instead of random in-context examples, we use the top-$k$ examples from the training set with the highest cosine similarity to the query (GPT3.5Retr). This few-shot retrieval setting differs from the semantic neighbor discussed above, in that we provide both the input and output of each instance rather than solely supplying target entities as additional input.
Fine Tuning We fine-tune T5 Raffel et al. (2020) (more recent models may be used too; see our biomedical experiment §4.1 fine-tuning an LLM). We observe that the generation models tend to copy phrases from the background context. For example, given the context “…hierarchical tables challenge numerical reasoning …”, the model will generate “hierarchical table reasoning for question answering” as the top prediction. For generating suggestions of novel ideas, we wish to discourage overly copying from the background context. We introduce a new in-context contrastive objective, where negative examples are taken from the text in the input (e.g., in Figure 2, the in-context negatives are plms, pre-training, etc). We compute an InfoNCE loss Oord et al. (2018) over the hidden states of the decoder, aiming to maximize the probability of the ground truth against those of in-context negatives:
3.3 Iterative Novelty Boosting with Retrieval
We further improve the novelty of generated ideas with a new iterative retrieve-compare-update scheme. Conceptually, we consider a novelty-inducing penalty $\gamma_{\text{nov}}(\mathcal{I},\mathcal{R})$ that penalizes ideas $\mathcal{I}$ that are too “close” to existing ideas in literature reference examples $\mathcal{R}$. $\gamma_{\text{nov}}(\mathcal{I},\mathcal{R})$ is included during in-context learning and inference, providing numerical feedback in the form of a score reflecting similarity to existing work. We wish to minimize this score while ensuring $\mathcal{I}$ remains relevant to the background context $\mathcal{B}$; we do so iteratively by (1) retrieving related work from $\mathcal{R}$, (2) measuring degree of novelty, (3) instructing the model to update $\mathcal{I}$ to be more novel w.r.t $\mathcal{R}$, conditioning on $\mathcal{B}$.
[…]
@@ 4 Experiments
4.1 Human Evaluation
We present four human evaluation studies, exploring different facets of our problem and approach.
Study I We recruit six volunteer NLP experts with graduate-level education to rate the system. Raters are told to envision an AI assistant that suggests new paper ideas. We randomly select 50 instances (background+seed) from the gold subset. Each annotator receives ten instances, each paired with system outputs from different model variants (Table 2). We ask raters to assess idea quality by considering each output’s relevance to the context, novelty, clarity, and whether the idea is reasonable (positive ratings are dubbed “helpful” as shorthand, indicating they pass the multiple considerations). We observe moderately high rater agreement. [^6] The agreement scores are in Table 13 Appendix C. Raters are blind to the condition, and system outputs are randomly shuffled across instances.
We instruct annotators to only provide positive ratings to ideas sufficiently different from the input context. In Study I, we ask raters not to anticipate groundbreaking novelty from the system but rather a narrower expectation of quality and utility; in Study II below, we enrich the analysis to examine ranking between top models and also “raise the bar” and compare to actual ideas from papers. [^7] Full evaluator guidelines are in Appendix C. The sample annotations are in Table 11.
In a preliminary experiment, we also collected human ratings for GPT4-ZS (zero-shot) vs. GPT4-FS (few-shot) using the same criteria, finding GPT4-FS ranked higher in 65% of cases, with the rest mostly tied; thus, zero-shot GPT-4 was left out of the remainder of study I and subsequent studies to reduce annotation effort and cost.
Study I: Results Overall, GPT4FS and GPT4FS+KG outperform other models by a wide margin (Table 2). Apart from GPT4, T5+SN+CL performs best compared to other baselines, given its stronger prior knowledge of useful background citations. In general, GPT3.5 models performed worse than fine-tuned T5 and its variants, which echoes results in other work in the scientific NLP domain Jimenez Gutierrez et al. (2022). GPT4 outputs tended to be longer, which may partially explain higher human preference.
@@ 4.1 Human Evaluation
Table 2: Percent (%) of total votes each system output receives from human raters for the idea sentence prediction task. H denotes a helpful output, while U denotes an unhelpful output. “3FS” refers to the GPT3.5FS. “3Rt” refers to the GPT3.5Retr. “4” refers to GPT4FS, and “4+KG” refers to the GPT4FS+KG. “T5+SN” refers to the T5+SN+CL. GPT4FS and GPT4FS+KG are rated much higher. While GPT4FS has a slightly higher rating than the KG variant, a further human study reveals that GPT4FS+KG often leads to more technical depth (§4.1).
@@
Study II: GPT4 Comparisons to Real Papers We conduct a follow-up human study of close competitors GPT4FS and GPT4FS+KG with a subset of the annotators. In this study, model outputs are now ranked, unlike the binary classification of helpful/not in Study I. Suggestions are ranked according to the level of technical detail and innovation in comparison to each other—i.e., ranking which of GPT4FS and GPT4FS+KG had a higher degree of technical detail and novelty, or whether they are roughly the same (tied). Finally, outputs are rated versus the ground truth idea, according to whether or not the suggestions were roughly at the same level of technical detail and innovation as the original paper’s idea, or significantly lower.
@@
Study II: Results Overall, GPT4FS+KG is found to have higher technical detail in 48% of the compared pairs, and found to be less incremental (more novel) in 45% of the pairs. Among the remaining 52%/55% (respectively), the vast majority are ties, indicating that whenever GPT4FS+KG is not favored, it is of roughly the same quality as GPT4FS, but not vice versa. However, the most crucial aspect is comparing the results against the original ground truth idea on the quality of innovation. Here, we find that in 85% of comparisons, the ground truth is considered to have significantly higher technical level and novelty; and in the remaining 15%, the ground truth was ambiguous or lacking additional context from the paper abstract. This points to a major challenge in obtaining high-quality idea generations using existing state-of-the-art models.
Study III: Iterative Novelty Boosting We ask five annotators to further compare the novelty-enhanced results against the initially generated ideas. We randomly select 70 instances (background+seed) from the sentence generation gold subset. We ask annotators to check whether the new ideas are different than the initial ideas (e.g., adding new information or approaches), and whether they are more novel (i.e., a new idea can be different, but not necessarily more novel). Since GPT4FS+SN outperforms other models, for this model we further instruct annotators to compare the novelty of the second iteration results against first iteration results.
Type GPT4FS +SN +CT +KG
1st Novelty $\Delta$ (%)
+54.4
+55.6
+47.8
+46.7
2nd Novelty $\Delta$(%)
+57.8
[…]
1st new terms $\Delta$
+23.1
+22.8
+22.1
+21.9
2nd new terms $\Delta$
+21.5
Table 3: Relative improvements of iterative novelty boosting. Iterations are applied to the ideas for which sufficiently similar related work is detected (§3.3). “1st Novelty” is % of the 1st iteration ideas that gained novelty over the initial idea, and “2nd Novelty” is the % of gain over the 1st iteration. Our method substantially increases novelty for ideas to which it is applied. To save annotation resources, we only annotate 2nd iteration results for the best-performing method (SN). We report the average number of new terms added, after filtering.
Study III: Results For SN, in the first iteration 88.9% of updated ideas are substantially different from initial ideas, and for 55.6% we are able to increase novelty/creativity (meaning that, e.g., if 100 examples were updated, we would gain 56 examples that are more novel). The 2nd iteration, further increases novelty for 57.8% of the ideas that continued to another iteration. For ideas not considered more novel after applying our method, we do not observe a drop in novelty—the method either increases or maintains novelty.
Ideas after novelty iterations are longer than initial ideas. We examine the new terms added after filtering 359 words, including stopwords, as we find many generic words and terms are often added (e.g., “novel model/method/approach”). While our method helps boost novelty, overall the model often tends to suggest combinations between popular concepts (§4.2). Novelty boosting seemed to often focus on adding dynamic/adaptive modeling, graph models and representations, the fusion of multiple modalities and sources—and sometimes all at once (e.g., “Dynamic Syntax-Aware Graph Fusion Networks (DSAGFN)”), and to explicitly compare against existing ideas from literature (Table 5).
Type Meditron +SN +CT +KG
Helpful(%)
35
80
60
50
Unhelpful(%)
65
20
40
50
vs. GT(%)
30
45
50
35
Table 4: Human evaluations results of each system output for the idea sentence prediction task on Biomedical Domain. “vs. GT” refers to percents which system outputs are better than ground truth ideas.
Type Content
Input Dong et al. (2022)
seed term: speech unit boundaries ; context (abridged): … generate partial sentence translation given a streaming speech input. existing approaches … break the acoustic units in speech, as boundaries between acoustic units in speech are not even. …
[…]
Initial idea
A pause prediction model to identify speech unit boundaries
[…]
Iteration 1
A method that leverages acoustic and linguistic features to predict speech unit boundaries dynamically, ensuring smooth transitions … differs from the existing research as it combines both acoustic properties and linguistic context … adapting to variations in speaker characteristics, speaking styles, and languages.
[…]
Iteration 2
A novel method called Adaptive Speech Unit Boundary Detection (ASUBD) … a combination of attention mechanisms to focus on relevant acoustic and linguistic features and reinforcement learning to guide the system to make optimal predictions of unit boundaries based on previous decisions…
[…]
Ground Truth
… an efficient monotonic segmentation module … accumulate acoustic information incrementally and detect proper speech unit boundaries.
Table 5: Example of iterative novelty iterations. Our novelty iteration method enhances ideas overall; however ideas are often based on superficial recombinations of common concepts, far from the technical depth of scientific papers.
@@
Domain Generalization Case Study Our framework is domain-agnostic and can be applied to other domains by changing the IE system used in the preprocessing procedure. To demonstrate this, we conduct an additional initial experiment in the biochemical domain. We follow a similar data creation procedure as for NLP papers. We collect a dataset from PubMed papers and use PubTator 3 Islamaj et al. (2021); Wei et al. (2022); Luo et al. (2023); Wei et al. (2023); Lai et al. (2023) as an IE system to extract a KG from paper abstracts. We use a sentence classifier trained on annotated abstracts Huang et al. (2020) to select background context. We fine-tune a state-of-the-art biomedical large language model Chen et al. (2023) on our data and evaluate on a test split past its pre-training cutoff date. [^8] More data and training details in Appendix A.2, B.2.3. We ask two biochemical domain experts with graduate-level education to evaluate the quality of the results as before, finding them to overall rate 80% of the generated directions positively. Finally, in contrast to NLP-domain experiments, evaluators were more satisfied with the generated outputs compared to the ground truth in terms of technical detail. Detailed results are in Table 4. However, this preliminary experiment was meant mainly to demonstrate the generality of our approach, and a more in-depth exploration of utility and quality is left for future work.
4.2 Error Analysis, Automated Metrics
Error Analysis Models often made generic suggestions, woven together with specific details copied directly from the context (e.g., “NLP with ML algorithms and sentiment analysis” for some problem X, or “data augmentation and transfer learning” for Y, or “BERT or RoBERTa” for Z). Our techniques reduced this behavior but did not fully solve it. GPT4 models, especially, seemed to generate generic descriptions of common steps in NLP workflows (e.g., “Data preprocessing: Clean the text data, remove unnecessary characters, perform tokenization…”). All models often copied and rephrased directly from the context. In certain cases, models applied simple logical modifications to the context; e.g., when contexts described problems such as “high latency” or “efficiency limitations”, the suggestions would include phrases such as “low latency” or “highly efficient”.
Automated Evaluation Analysis In open-ended tasks such as ours, automatic evaluations comparing system output to ground truth texts may be limited. Nonetheless, automated metrics such as ROUGE Lin (2004), BERTScore Zhang* et al. (2020) and BARTScore Yuan et al. (2021), that check the similarity between ground truth and generated output, may surface interesting findings. We find GPT-based models to be outperformed by T5-based models; GPT4 outputs are much longer than T5, explaining why they underperform in automatic metrics but outperform in human evaluations (§4.1). Generated sentences often follow certain templates (e.g., “In this paper, we propose a new … for …”), which also helps explain why T5 fine-tuned on many examples scores higher superficially. At the same time, our in-context contrastive examples which encourage novelty with respect to background context, helped models perform better than baseline fine-tuning by reducing reliance on copying. See results in Table 9 (Appendix B.3).